A student bikes to school by traveling first d_N = 0.900rm {miles} north, then d_W = 0.600rm {miles} west, and finally d_S = 0.100rm {miles} south.
Answer #1
If a bird were to start out from the origin (where the student starts) and fly directly (in a straight line) to the school, what distance d_b would the bird cover?
answered by: UCI Student
Answer #2
north .9 - .1 = .8
west .6

right triangle beacuse north is perpendicular to west

c^2 = a^2 + b^2
c^2 = .8^2 + .6^2

3,4 ,5 triangle = 2 (.3 , .4 , .5)
so answer is 1

or do it out
.8^2 = .64
.6^2 = .36
sum is 1
sqrt (1) = 1
answered by: Lady
Answer #3
[REMOVED]